### v1

# DogeRM: Equipping Reward Models with Domain Knowledge through Model Merging

Abstract
Reinforcement learning from human feedback (RLHF) is a popular strategy for aligning large language models (LLMs) with desired behaviors. Reward modeling is a crucial step in RLHF. However, collecting paired preference data for training reward models is often costly and time-consuming, especially for domain-specific preferences requiring expert annotation. To address this challenge, we propose the Domain knowledge merged Reward Model (DogeRM), a novel framework that integrates domain-specific knowledge into a general reward model by model merging. The experiments demonstrate that DogeRM enhances performance across different benchmarks and provide a detailed analysis showcasing the effects of model merging, showing the great potential of facilitating model alignment.
DogeRM: Equipping Reward Models with Domain Knowledge
through Model Merging
Tzu-Han Lin  Chen-An Li  Hung-yi Lee  Yun-Nung Chen National Taiwan University, Taipei, Taiwan {r12944034,b08902123}@csie.ntu.edu.twhungyilee@ntu.edu.twy.v.chen@ieee.org
1 Introduction
Modern large language models (LLMs), such as GPT-4 Achiam et al. (2023) and Gemini Team et al. (2023), showcase impressive capabilities across various tasks Gao et al. (2023); Beeching et al. (2023), with aligning their behavior with human preferences. Reinforcement learning from human feedback (RLHF) is a prominent technique for enhancing the alignment of desired behaviors in LLMs Christiano et al. (2017); Ziegler et al. (2019); Ouyang et al. (2022). A key component of RLHF is its reward models (RMs), which assess entire sentences generated by policy models. The reward signals produced by these RMs are instrumental in adjusting the parameters of the policy models, thus directly impacting the policy models’ effectiveness.
RMs are developed by training LLMs on paired preference data to simulate human judgment Ouyang et al. (2022). This preference data consists of two responses to a given user input, accompanied by a human-assigned label indicating which response is more preferred. However, gathering such preference data can be costly and time-consuming due to the requirement of human annotation Stiennon et al. (2020a). This challenge becomes more pronounced when handling domain-specific preference data, as it necessitates expertise from domain specialists.
Figure 1: The framework of DogeRM, where we merge the general RM with a domain-specific LM to create the domain-specific RM.
Recent developments have demonstrated the effectiveness of model merging techniques in strategically integrating multiple domain-specific models into a multi-domain model without requiring additional training Wortsman et al. (2022); Ilharco et al. (2023). Furthermore, domain-specific SFT data is relatively more accessible compared to preference data. Moreover, many high-quality domain-specific models are available on open-source platforms Wolf et al. (2019), which can be directly employed in the merging process. This brings us to consider a novel approach: Is it possible to equip reward models with domain knowledge through merging with domain-specific language models?
In this work, we propose Domain knowledge merged Reward Model (DogeRM), exploring the potential of merging a reward model trained on a general open-sourced preference dataset with a language model fine-tuned on domain-specific datasets, such as math and coding. An illustration of DogeRM is presented in Figure 1. We evaluate DogeRM using RewardBench Lambert et al. (2024), Auto-J Eval Li et al. (2024) and Best-of-N sampling on GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021). Our results demonstrate that DogeRM improves performance and can be generalized to different model architectures. We also conduct a comprehensive analysis to demonstrate the impact of model merging.
2 Related Work
Reward Modeling
RMs are crucial for aligning language models with human preferences, providing proxy rewards as training signals for policy models. Previous work has employed RL algorithms with RMs to guide language models towards human preferences in various NLP tasks Ziegler et al. (2020); Stiennon et al. (2020b); Wu et al. (2021); Nakano et al. (2022); Menick et al. (2022) and instruction-following Ouyang et al. (2022); Ramamurthy et al. (2023). In RLHF literature, RMs evaluate the quality of instructions and responses based on criteria like helpfulness and harmlessness Bai et al. (2022) or more fine-grained objectives Wu et al. (2023).
Several open-source paired preference datasets are available for training RMs for RLHF, such as OpenAI Summarization Stiennon et al. (2020b), HH-RLHF Bai et al. (2022), SHP Ethayarajh et al. (2022), Ultrafeedback Cui et al. (2023), PKU-SafeRLHF Ji et al. (2023), HelpSteer Wang et al. (2023), Nectar Zhu et al. (2023), and UltraInteract Yuan et al. (2024). However, most datasets are not domain-specific. To address this, our work focuses on merging RMs with domain-specific language models, aiming to equip RMs with domain knowledge.
Model Merging
Model merging integrates multiple task-specific models into a single unified model without additional training. A straightforward approach involves averaging parameters from models fine-tuned from the same initial model Wortsman et al. (2022). Another method employs weighted averaging of model parameters Matena and Raffel (2022); Jin et al. (2023).
Another innovative approach involves creating task vectors by subtracting the weights of a pre-trained model from those of the same model after fine-tuning for a specific task. This method showcases the flexibility and composability of these vectors through arithmetic operations Ilharco et al. (2023); Yadav et al. (2024); Huang et al. (2024).
Some recent work focused on model merging to align with user preferences. They interpolated model parameters fine-tuned on diverse rewards Rame et al. (2024a); Jang et al. (2023); Wang et al. (2024a), or merging RMs for combining different aspects of rewards Rame et al. (2024b).
However, these approaches still require domain-specific preference data to integrate domain knowledge. Our work eliminates the requirement of domain-specific preference data and focuses on incorporating domain-specific knowledge into RMs through model merging.
3 Methodology
3.1 Reward Modeling
To train a reward model, we replace the decoding layer of a transformer-based pre-trained language model with a linear regression layer. This new layer projects the logits from the final transformer layer to a scalar, representing the reward of a given input.
Given an input prompt $x$, the chosen response $y_{c}$, and the rejected response $y_{r}$, we use the following loss function to optimize our reward model:
$\mathcal{L}_{\text{RM}}=-\log{[\sigma(r(x,y_{c}))-\sigma(r(x,y_{r}))]}\vspace{-5pt}$ (1)
where $r(x,y_{c})$ is the reward of chosen response, $r(x,y_{r})$ is the reward of rejected response, and $\sigma(\cdot)$ is the logistic function.
Model Reward Bench Auto-J Eval
Chat Chat-Hard Safety Reasoning Code Math Others
Code Math
LLaMA-2 RM 95.8 47.6 44.6 78.9 68.2 76.2 84.4 79.2
Ours (+ MetaMath) 95.8 44.5 43.5 85.7 79.6 79.8 87.5 79.3
Ours (+ MAmmoTH) 96.1 44.7 43.8 84.1 85.2 79.8 87.5 79.7
Ours (+ Code Model) 96.1 45.6 43.9 84.3 71.8 82.1 87.5 79.7
Table 1: Performance comparison on various benchmarks.
3.2 Model Merging
Our proposed method merges the parameters of a supervised fine-tuned language model, denoted as $\theta^{\text{SFT}}$, with those of a reward model, $\theta^{\text{RM}}$, both initialized from the same pre-trained model $\theta$.
We divide $\theta^{\text{SFT}}$ into three disjoint parts:
$\theta^{\text{SFT}}=\{\theta^{\text{SFT}}_{\text{emb}},\theta^{\text{SFT}}_{\text{trans}},\theta^{\text{SFT}}_{\text{dec}}\}\vspace{-5pt}$ (2)
where $\theta^{\text{SFT}}_{\text{emb}},\theta^{\text{SFT}}_{\text{trans}},\theta^{\text{SFT}}_{\text{dec}}$ represent the embedding, transformer, and decoding layers’ parameters, respectively.
Similarly, we also divide $\theta^{\text{RM}}$ into three parts:
$\theta^{\text{RM}}=\{\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{reg}}\}\vspace{-5pt}$ (3)
where $\theta^{\text{RM}}_{\text{emb}},\theta^{\text{RM}}_{\text{trans}},\theta^{\text{RM}}_{\text{reg}}$ denote the parameters for the embedding, transformer, and regression layer, respectively.
For embedding layer parameters, we apply a weighted average to common token embeddings:
$\theta^{\text{MERGE}}_{\text{emb},t_{i}}=\lambda\cdot\theta^{\text{SFT}}_{\text{emb},t_{i}}+(1-\lambda)\cdot\theta^{\text{RM}}_{\text{emb},t_{i}}$ (4)
where $t_{i}$ is a common token to both models, $\theta_{\text{emb},t_{i}}$ is the corresponding embedding, and $\lambda$ are weights for SFT parameters ranging from 0 to 1.
As for the unshared tokens, we directly use the embedding from their corresponding source model.
$\theta^{\text{MERGE}}_{\text{emb},t_{i}}=\begin{cases}\theta^{\text{SFT}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to SFT model}\\ \theta^{\text{RM}}_{\text{emb},t_{i}}&\text{If $t_{i}$ is unique to RM}\end{cases}$ (5)
For the transformer layers, we perform a weighted average directly since both models are initialized from the same pre-trained model:
$\theta^{\text{MERGE}}_{\text{trans}}=\lambda\cdot\theta^{\text{SFT}}_{\text{trans}}+(1-\lambda)\cdot\theta^{\text{RM}}_{\text{trans}}$ (6)
Finally, we derive the merged reward model $\theta^{\text{MERGE}}$ by combining $\theta^{\text{MERGE}}_{\text{emb}}$, $\theta^{\text{MERGE}}_{\text{trans}}$, and the reward model’s regression layer $\theta^{\text{RM}}_{\text{reg}}$:
$\theta^{\text{MERGE}}=\{\theta^{\text{MERGE}}_{\text{emb}},\theta^{\text{MERGE}}_{\text{trans}},\theta^{\text{RM}}_{\text{reg}}\}$ (7)
4 Experiments
4.1 Experimental Setup
Reward Model
To fine-tune the backbone of our reward model, we utilize the 10k SFT split from Alpacafarm Dubois et al. (2023). For reward modeling, we employ the UltraFeedback Cui et al. (2023). The details of training and these datasets are presented in Appendix A and C, respectively.
Domain-Specific SFT
For the math LLMs, we adopt the open-source models, MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a), both of which are fine-tuned from LLaMA-2-7B. For code generation LLM, since we could not find open-source models with detailed training information, we use OSS-Instruct and Magicoder-Evol-Instruct Wei et al. (2024) to fine-tune LLaMA-2-7B ourselves. We refer to this model as the Code Model in the following sections. The details of code fine-tuning datasets and math models are presented in Appendix C and D, respectively.
4.2 Evaluation
We evaluate the reward models using two benchmarks, RewardBench Lambert et al. (2024) and Auto-J Eval Li et al. (2024). These benchmarks provide paired instruction-completion data, with the preferred completion annotated as chosen and the other as rejected, using accuracy as the evaluation metric. We use the core set of RewardBench, focusing primarily on the reasoning category to evaluate the model’s abilities in code and mathematical reasoning. For Auto-J Eval, we use pair-wise testing data and categorize the dataset into three categories: code, math, and others, following Yuan et al. 2024. Additionally, to further test our reward model’s effectiveness in enhancing model performance through reranking, we conduct best-of-N sampling on zero-shot prompted responses from llama-2-7B-chat for GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021). We present results using a weight factor of $\lambda=0.35$ for the main findings, with an analysis of the impact of $\lambda$ detailed in section 4.4 and results for different values of $\lambda$ presented in Appendix F.
4.3 Results
RM Benchmarks
The main results RM benchmark are shown in Table 1. Merging the LLaMA-2 RM with MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a) improves math performance on RewardBench by 11.4% and 17%, respectively, and coding performance by 5.2% and 5.8%, respectively. Similar enhancements are seen on Auto-J Eval, with gains in both math and coding. Merging our LLaMA-2 RM with the Code Model further improves coding performance on RewardBench and Auto-J Eval by 5.4% and 6%, respectively, along with noticeable improvements in math performance on both benchmarks. Although our methods improve the reasoning domain, performance in other domains did not significantly degrade.
Best-of-N Sampling
Figure 2 shows the results, with accuracy improvements on GSM8K and MBPP. At the best-of-16 setting, merging with math models improves GSM8K by 5%, and merging with the Code Model enhances MBPP by 1.4%.
(a) + MetaMath/MAmmoTH on GSM8K.
(b) + Code Model on MBPP.
Figure 2: Best-of-N results. Merging with domain-specific models improves reranking accuracy. Topline: Pass@N, the probability of obtaining at least one correct solution out of N responses. Baseline: LLaMA-2 RM.
4.4 Analysis
Effect of Weight Factor $\lambda$
To further investigate how the weight factor $\lambda$ affects our method’s performance, we test various values of $\lambda$ ranging from 0 to 1 in increments of 0.05, observing the performance changes across these values on RewardBench. Figure 3 shows that performance degrades when $\lambda$ is large. We suggest setting $\lambda$ between 0.2 and 0.5 to achieve better results.
Reward Differences
We delve deeper into how model merging affects the output of reward models by examining the value of the reward signals corresponding to chosen and rejected prompts in RewardBench. Figure 3 illustrates the distribution before and after merging. In the math subset, we notice that the difference in reward scores between the chosen and rejected prompts initially increases and then decreases as $\lambda$ varies from 0 to 1. Conversely, in the code subset, this difference consistently decreases. We hypothesize that this discrepancy arises because the original reward model inherently excels in the code subset.
(a) + MAmmoTH on RewardBench math subset.
(b) + Code Model on RewardBench code subset.
(c) + MAmmoTH on RewardBench math subset.
(d) + Code Model on RewardBench code subset.
Figure 3: The impact of different value of $\lambda$ on RewardBench math and code subsets. (a)(b): Accuracy; (c)(d): Reward difference between chosen and rejected prompts.
Generalizability
To test the adaptability of DogeRM to different model architectures, we use an open-source Mistral-based Jiang et al. (2023) RM Ray2333 (2024) merging with Misrtral-based MAmmoTH2-7B-Plus Yue et al. (2024b). Details of these models are presented in Appendix D. The results for $\lambda=0.35$ in reasoning domains on RM benchmarks and best-of-N sampling on GSM8K, with N=16, are presented in Table 2. Our method improves math performance by 30% on RewardBench and 3% on Auto-J Eval. Additionally, we enhance reranking performance on GSM8K by 2.4%. These results demonstrate the adaptability of our methods to different model architectures. The results for different $\lambda$ are presented in Appendix F.
Model Reward Bench Auto-J Eval Best-of-16
Code Math Code Math GSM8K
Mistral RM 93.5 55.0 88.1 87.5 44.2
+ MAmmoTH2-Plus 92.6 85.0 88.1 90.6 46.6
Table 2: Performance of Mistral-based models on various benchmarks and best-of-16 results. Our methods show improvements across RM benchmarks and in best-of-16 sampling on GSM8K.
5 Conclusion
In this work, we introduce a novel approach, DogeRM, which integrates domain knowledge into RM by merging it with the SFT model. We demonstrate that DogeRM enhances performance on math and coding benchmarks and can be generalized to different architectures. A series of analyses show that DogeRM effectively affects the reward signal corresponding to chosen and rejected prompts.
Limitations
There are several limitations in our work: (1) We only tested our framework in the math and coding domains; other domains such as medicine, finance, and law have not been explored. (2) Our method was tested exclusively on 7B models, and we have not evaluated its performance on models of larger or smaller sizes. (3) While our framework is compatible with various merging techniques, such as TIES-Merge Yadav et al. (2024), we did not investigate the actual effects of these different methods.
Ethics Statement
While our method effectively equips reward models with domain knowledge, it does not eliminate the inherent biases within these models. Further investigation is needed to explore the impact of these inherited biases in the original reward models.
Acknowledgments
We thank the National Center for High-performance Computing (NCHC) for providing computational and storage resources. We also thank Yen-Ting Lin and Chao-Wei Huang for the valuable discussions.
Appendix A Training Details
We use V100 GPUs for training models. We spent 2 hours training the backbone of our LLaMA-2 RM, 8 hours training our LLaMA-2 RM, and 12 hours training our Code Model. Since V100 did not support bf16, we adopted mixed precision training (fp16) for both SFT and Reward Modeling.
A.1 Supervised Fine-Tuning (SFT)
We use LlamaFactory Zheng et al. (2024) for supervised fine-tuning (SFT). For fine-tuning the backbone of our LLaMA-2 RM, we use Alpacafarm Dubois et al. (2023) with a learning rate of 1e-5 and a batch size of 128. For code generation, we follow a training procedure similar to Wei et al. (2024). First, we use OSS-Instruct to fine-tune LLaMA-2-7B Touvron et al. (2023) for 2 epochs. Then, we continuously fine-tune the model with Magicoder-Evol-Instruct for 1 epoch. The learning rate for both stages is 1e-5, and the effective batch size is 128.
A.2 Reward Modeling
For reward modeling, we modify the sample code provided by TRL von Werra et al. (2020). We trained the backbone model described in the previous section on UltraFeedback Cui et al. (2023) for 1 epoch, using a learning rate of 1e-5 and a batch size of 32.
Appendix B Prompt Template
For LLaMA-2 based models, we use the same prompt template as LLaMA-2-Chat model, as shown below:
<s>[INST] <<SYS>>
{System Prompt}
<</SYS>>
[…]
We use this template for both SFT and reward modeling. For Mistral-based models, the prompt template is modified by removing the system prompt part:
[…]
The default system prompt we used in SFT and reward modeling aligns with the original system prompt for LLaMA-2-Chat model:
You are a helpful, respectful and honest assistant. Always
answer as helpfully as possible, while being safe. Your
answers should not include any harmful, unethical, racist,
sexist, toxic, dangerous, or illegal content.
Please ensure that your responses are socially unbiased
and positive in nature. If a question does not make any
sense, or is not factually coherent, explain why instead
of answering something not correct. If you don’t know the
answer to a question, please don’t share false information.
The system prompt used in prompting LLaMA-2-7B-Chat for Best-of-N sampling on GSM8K is:
You are a math problem solver. Please think step by step and
demonstrate your calculation steps. After your reasoning
steps, you should generate the answer by following the for-
mat starting with ’The answer is’
The system prompt used in prompting LLaMA-2-7B-Chat for Best-of-N sampling on MBPP is:
    Write Python code to solve the task.
Appendix C Dataset Details
Alpacafarm Dubois et al. (2023)
The Alpacafarm dataset consists of 52k instructions as well as response generated by text-davinci-003 model OpenAI (2023) from the original Alpaca dataset Taori et al. (2023). Alpacafarm splits the datasets into 10k ’sft’ subset for instruction fine-tuning, 10k ’pref’ subset for preference learning, 20k ’unlabeled subset for training such as PPO, and 2k ’val’ subset for validation. We only utilize the 10k ’sft’ subset for fine-tuning the backbone of our reward model.
UltraFeedback Cui et al. (2023)
This dataset consists of 64k prompts from sources including UltraChat Ding et al. (2023), ShareGPT Chiang et al. (2023), Evol-Instruct Xu et al. (2024), TruthfulQA Lin et al. (2022), FalseQA Hu et al. (2023), and FLAN Longpre et al. (2023). The responses are generated by a pool of different LLMs. The preferences are generated by GPT-4 Achiam et al. (2023). In our experiment, we use a cleaned version of UltraFeedback [^1] https://huggingface.co/datasets/argilla/ultrafeedback-binarized-preferences-cleaned Bartolome et al. (2023), which removes TruthfulQA contamination and uses the average of the preference ratings.
OSS-Instruct & Magicoder Evol-Instruct Wei et al. (2024)
OSS-Instruct consists of 75k synthesized data collected by prompting ChatGPT Achiam et al. (2023) to generate a coding problem and solution based on a seed code snippet from an open-sourced platform. The Magicoder Evol-Instruct dataset, based on the work in Luo et al. (2023), uses an open-source implementation theblackcat102 (2023) that has been further decontaminated, resulting in 110k data points for fine-tuning. Both OSS-Instruct and Magicoder Evol-Instruct are used to fine-tune the Code Model for merging.
RewardBench Lambert et al. (2024)
RewardBench is a benchmark designed to evaluate reward models (RMs). The datasets are categorized into core sets and prior sets. The prior sets consist of testing sets from open-sourced preference dataset such as OpenAI Summarization Stiennon et al. (2020b), Anthropic Helpful split Bai et al. (2022), Anthropic HHH subset Askell et al. (2021), and Stanford Human Preference (SHP) Ethayarajh et al. (2022).
We utilize the core sets for evaluation, which include four categories: chat, chat-hard, reasoning, and safety. The chat category collects data from AlpacaEval Li et al. (2023) and MT Bench Zheng et al. (2023) to assess RMs’ basic ability to discern correct responses in open-ended dialogue. Chat-Hard incorporates data from MT Bench Zheng et al. (2023) with similar ratings and LLMBar Zeng et al. (2024) data designed to challenge LLM-based judges. The reasoning category includes math data selected from PRM800K Lightman et al. (2024), where the prompt is the reference answer and the rejected prompt is a wrong solution generated by GPT-4 Achiam et al. (2023). The coding data utilizes HumanEvalPack Muennighoff et al. (2024), augmenting HumanEval Chen et al. (2021) across six programming languages, with the prompt being the reference solution and the rejected prompt being buggy solutions. Safety category comprises data from XSTest Röttger et al. (2024), Do-Not-Answer Wang et al. (2024b), and an in-development refusals dataset at AI2, aiming to accurately test models’ ability to refuse dangerous content and avoid incorrect refusals triggered by similar words.
Auto-J Eval Li et al. (2024)
Auto-J Eval’s pairwise testing set includes examples from various sources: OpenAI Summarization Stiennon et al. (2020a), WebGPT Nakano et al. (2022), Stanford SHP Ethayarajh et al. (2022), Synthetic GPT-J Havrilla (2023), and PKU-SafeAlignment Ji et al. (2023). GPT-4 Achiam et al. (2023) serves as the annotator. The dataset consists of categories including Summarization, Exam Questions, Code, Creative Writing, Functional Writing, Rewriting, General Communication, and NLP Tasks. We exclude the tied examples and re-group the data into Code, Math (extract from Exam Questions category), and Others, following Yuan et al. 2024.
GSM8K Cobbe et al. (2021)
This dataset consists of 8.5K grade school-level math problems. We use the prompt from the testing set to perform Best-of-N sampling in a zero-shot manner.
MBPP Austin et al. (2021)
This dataset consists of 1,000 crowd-sourced Python programming problems, which are entry-level problems covering standard libraries, programming, and so on. We use the testing set to perform Best-of-N sampling in a zero-shot manner.
Appendix D Open-Source Model Details
MetaMath Yu et al. (2024)
We use the LLaMA-2-7B based model fine-tuned by the authors for merging. The MetaMath-7B models are trained on the MetaMathQA dataset, which the authors curated by bootstrapping problems from GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021).
MAmmoTH Yue et al. (2024a)
We merge our RM with the MAmmoTH-7B model, a LLaMA-2-7B based model fine-tuned on the MathInstruct dataset. This dataset combines a diverse range of math problems and hybrid rationales curated by the author.
Mistral-RM Ray2333 (2024)
We utilize a Mistral-based reward model trained on a variety of preference datasets to test our framework’s adaptability. This reward model is initialized from Mistral-7B-Instruct-v0.2. Detailed information about the training setup can be found in the author’s blog [^2] https://www.notion.so/Reward-Modeling-for-RLHF-abe03f9afdac42b9a5bee746844518d0.
MAmmoTH2-Plus Yue et al. (2024b)
To test the adaptability of our framework across different model architectures, we use the MAmmoTH2-7B-Plus and merge it with the Mistral RM. This model is fine-tuned from the MAmmoTH2-7B, which is fine-tuned from Mistral-7B-Instruct-v0.2, on public instruction tuning datasets to further enhance performance.
Appendix E Case Study
We conducted a case study on the predictions of the reward model. We examined the changes in chosen/rejected prompt rewards on the coding subset of RewardBench before and after merging the reward models. We identified that the problems initially predicted incorrectly but corrected after merging can be categorized into math, string operations, and bit operations. Of the four problems correctly predicted before merging but incorrectly predicted after merging, two were related to list/sorting operations, while the others involved mathematical operations such as modular arithmetic and derivatives.
E.1 Passed Case after Merging (Math)
Instruction
Write a Python function ‘max_fill(grid, capacity):
import math‘ to solve the following problem:
You are given a rectangular grid of wells. Each row represents a single well,
and each 1 in a row represents a single unit of water.
Each well has a corresponding bucket that can be used to extract water from it,
and all buckets have the same capacity.
Your task is to use the buckets to empty the wells.
Output the number of times you need to lower the buckets.
Example 1:
Input:
grid : [[0,0,1,0], [0,1,0,0], [1,1,1,1]]
bucket_capacity: 1
Output: 6
Example 2:
Input:
grid : [[0,0,1,1], [0,0,0,0], [1,1,1,1], [0,1,1,1]]
bucket_capacity : 2
Output: 5
Example 3:
Input:
grid: [[0,0,0], [0,0,0]]
bucket_capacity : 5
Output: 0
Constraints:
* all wells have the same length
* 1 <= grid.length <= 10^2
* 1 <= grid[:,1].length <= 10^2
* grid[i][j] -> 0 | 1
* 1 <= capacity <= 10
Chosen Response
return sum([math.ceil(sum(arr)/capacity) for arr in grid])
Rejected Response
return sum([math.floor(sum(arr)/capacity) for arr in grid])
E.2 Passed Case after Merging (String)
Instruction
Write a Python function ‘concatenate(strings: List[str]) -> str‘ to solve the following problem:
Concatenate list of strings into a single string
$>$ concatenate([])
$>$ concatenate([’a’, ’b’, ’c’])
’abc’
Chosen Response
return ’’.join(strings)
Rejected Response
return ’ ’.join(strings)
E.3 Passed Case after Merging (Bit Operation)
Instruction
Write a Python function ‘string_xor(a: str, b: str) -> str‘ to solve the following problem:
Input are two strings a and b consisting only of 1s and 0s.
Perform binary XOR on these inputs and return result also as a string.
$>$ string_xor(’010’, ’110’)
’100’
Chosen Response
def xor(i, j):
    if i == j:
        return ’0’
    else:
        return ’1’
    return ’’.join(xor(x, y) for x, y in
zip(a, b))
Rejected Response
def xor(i, j):
    if i == j:
        return ’1’
    else:
        return ’0’
    return ’’.join(xor(x, y) for x, y in
zip(a, b))
E.4 Failed Case after Merging (List/Sorting)
Instruction
Write a Python function ‘sort_array(array)‘ to solve the following problem:
Given an array of non-negative integers, return a copy of the given array after sorting,
you will sort the given array in ascending order if the sum( first index value, last index value) is odd,
or sort it in descending order if the sum( first index value, last index value) is even.
Note:
* don’t change the given array.
Examples:
* sort_array([]) => []
* sort_array([5]) => [5]
* sort_array([2, 4, 3, 0, 1, 5]) => [0, 1, 2, 3, 4, 5]
* sort_array([2, 4, 3, 0, 1, 5, 6]) => [6, 5, 4, 3, 2, 1, 0]
Chosen Response
return [] if len(array) == 0 else sorted(array, reverse= (array[0]+array[-1]) % 2 == 0)
Rejected Response
return [] if len(array) == 0 else sorted(array, reverse= (array[0]+array[-1]) % 2 != 0)
E.5 Failed Case after Merging (Math)
Instruction
Write a Python function ‘derivative(xs: list)‘ to solve the following problem:
xs represent coefficients of a polynomial.
xs[0] + xs[1] * x + xs[2] * x^2 + ....
Return derivative of this polynomial in the same form.
$>$ derivative([3, 1, 2, 4, 5])
[1, 4, 12, 20]
$>$ derivative([1, 2, 3])
[2, 6]
Chosen Response
return [(i * x) for i, x in enumerate(xs)][1:]
Rejected Response
return [(i * x) for i, x in enumerate(xs)]
Appendix F Full Results
Full results with different values of $\lambda$ on Best-of-N sampling and RM benchmarks are presented here.
F.1 Best-of-N
Figure 4 and  5 demonstrate the results of Best-of-N sampling on GSM8K when merging our LLaMA-2 RM with MetaMath-7B Yu et al. (2024) and MAmmoTH-7B Yue et al. (2024a), respectively. DogeRM shows consistent improvement across different models being merged.
Figure 6 shows the result of Best-of-N sampling on MBPP when merging our LLaMA-2 RM with the Code Model. DogeRM improves the reranking accuracy, indicating that our method can be generalized to different tasks.
Finally, Figure 7 shows the results when merging the Mistral RM Ray2333 (2024) with MAmmoTH2-7B-Plus Yue et al. (2024b). DogeRM improves the reranking accuracy at an N=16 setting by 2.88%, indicating that our method can be generalized to different model architectures.
F.2 RewardBench
Figure 8 and 9 shows the results on different categories. We further split the reasoning category into math and coding. Merging LLaMA-2 RM with math models shows consistent improvement in both Math and Coding. The performance drop in chat-hard and safety categories can be observed.
Figure 10 shows the result of merging LLaMA-2 RM with the Code Model. We observe improvements in both the Math and Coding, with a performance drop in both the chat-hard and safety categories.
Finally, Figure 11 shows the result of merging Mistral RM with MAmmoTH2-7B-Plus. We improve accuracy on the math subset by 30%, while the improvement on the coding subset is minor, likely because the original RM already achieved high accuracy on this subset. An improvement in the chat-hard category can also be observed, contrary to previous cases, but a performance degradation in the safety category is found.
We believe that the performance degradation in safety aligns with observations from Yuan et al. 2024, which indicate that removing safety data from the RM training set improves reasoning performance, suggesting that modeling safety may hurt reasoning. As for the chat-hard category, we did not observe consistent performance degradation across all combinations. A deeper investigation into this is left for future work. Despite these issues, our method can effectively equip the LLaMA-2 RM with domain-specific knowledge, a finding that holds across different domains as well as different model architectures.
F.3 Auto-J Eval
The results of merging LLaMA-2 RM with math models are presented in Figure 12 and 13, showing improvements in both the Code and Math subsets. A similar observation can be found in Figure 14, which shows the result of merging LLaMA-2 RM with the Code Model, and Figure 15, which shows the result of merging Mistral RM with MAmmoTH-2-7B-Plus. These results support the conclusion that DogeRM can equip RMs with domain-specific knowledge.
(a) Best-of-2
(b) Best-of-4
(c) Best-of-8
(d) Best-of-16
Figure 4: Full results of LLaMA-2 RM + MetaMath on GSM8K.
(a) Best-of-2
(b) Best-of-4
(c) Best-of-8
(d) Best-of-16
Figure 5: Full results of LLaMA-2 RM + MAmmoTH on GSM8K.
(a) Best-of-2
(b) Best-of-4
(c) Best-of-8
(d) Best-of-16
Figure 6: Full results of LLaMA-2 RM + Code Model on MBPP.
(a) Best-of-2
(b) Best-of-4
(c) Best-of-8
(d) Best-of-16
Figure 7: Full results of Mistral RM + MAmmoTH2-Plus on GSM8K.
(a) Chat
(b) Chat-Hard
(c) Safety
(d) Code
(e) Math
Figure 8: Full results of LLaMA-2 RM + MetaMath on Reward Bench.
(a) Chat
(b) Chat-Hard
(c) Safety
(d) Code
(e) Math
Figure 9: Full results of LLaMA-2 RM + MAmmoTH on Reward Bench.
(a) Chat
(b) Chat-Hard
(c) Safety
(d) Code
(e) Math
Figure 10: Full results of LLaMA-2 RM + Code Model on Reward Bench.
(a) Chat
(b) Chat-Hard
(c) Safety
(d) Code
(e) Math
Figure 11: Full results of Mistral RM + MAmmoTH2-Plus on Reward Bench.
(a) Code
(b) Math
(c) Others
Figure 12: Full results of LLaMA-2 RM + MetaMath on Auto-J Eval.
(a) Code
(b) Math
(c) Others
Figure 13: Full results of LLaMA-2 RM + MAmmoTH on Auto-J Eval.
(a) Code
(b) Math
(c) Others
Figure 14: Full results of LLaMA-2 RM + Code Model on Auto-J Eval.
(a) Code
(b) Math
(c) Others
Figure 15: Full results of Mistral RM + MAmmoTH2-Plus on Auto-J Eval.